OpenStudy (anonymous):

Karen's penny back is 7/10 full. After she removes 200 pennies, it is 1/2 full. How many pennies can Karen's bank hold?

OpenStudy (sleepyhead314):

Hello @brittanyf188 and Welcome to OpenStudy! :) if Karen removed 200 pennies and the fraction goes from 7/10 to 5/10 7/10 - 5/10 = 2/10 this means that 200 pennies = 2/10 of the maximum number of pennies karen's bank can hold... 2/10 + 2/10 + 2/10 + 2/10 + 2/10 = 10/10 = full so this means that we should do 200 + 200 + 200 + 200 + 200 = [answer] = full

OpenStudy (mathmale):

This reasoning sounds correct. Anyone willing and able to check it out? How many coins can the bank hold, and after 200 coins are removed (from the original 700 in the bank), will the result be 500, or half of the number of coins the bank can hold?

OpenStudy (mathmale):

Alternative approach: Consider solving the following for x: (7/10)x-200=(5/10)x